### v1

# Few shot clinical entity recognition in three languages: Masked language models outperform LLM prompting

Abstract
Large Language Models are becoming the go-to solution for many natural language processing tasks, including in specialized domains where their few-shot capacities are expected to yield high performance in low-resource settings. Herein, we aim to assess the performance of Large Language Models for few shot clinical entity recognition in multiple languages. We evaluate named entity recognition in English, French and Spanish using 8 in-domain (clinical) and 6 out-domain gold standard corpora. We assess the performance of 10 auto-regressive language models using prompting and 16 masked language models used for text encoding in a biLSTM-CRF supervised tagger. We create a few-shot set-up by limiting the amount of annotated data available to 100 sentences. Our experiments show that although larger prompt-based models tend to achieve competitive F-measure for named entity recognition outside the clinical domain, this level of performance does not carry over to the clinical domain where lighter supervised taggers relying on masked language models perform better, even with the performance drop incurred from the few-shot set-up. In all experiments, the CO2 impact of masked language models is inferior to that of auto-regressive models. Results are consistent over the three languages and suggest that few-shot learning using Large language models is not production ready for named entity recognition in the clinical domain. Instead, models could be used for speeding-up the production of gold standard annotated data.
keywords:
Deep Learning, Natural Language Processing, Named Entity Recognition, Few-shot Learning, Large Language Models
[…]
1 INTRODUCTION
Electronic Health Records (EHR) are rich sources of clinical information fushman2009decision , which often appear in unstructured text only escudie2017novel . Efficiently extracting information from EHRs into a more structured form can help advance clinical research, public health surveillance and automatic clinical decision support wang2018applications .
Named Entity Recognition (NER) is a critical primary step in information extraction. It consists in identifying mentions of relevant entities in text. In the context of clinical information extraction from EHRs, these can be mentions of clinical entities such as disorders or drugs. Extracting these entities can particularly benefit concept normalization cho2017normalization ; wajsburt2021normalization ; sung2022normalization as well as facilitate interpreting patient profiling and phenotyping gerardin2022multilabel . While general-domain NER (identifying entities such as persons and locations) has received much attention in the Natural Language Processing (NLP) community, clinical NER is widely considered as a harder problem : clinical entities are often jargon or ambiguous, and clinical texts have a nonstandard phrasal structure luo2020n2c2 ; leaman2015challenges .
Language Models have progressively become the main approach for tackling NER li2022surveyner ; wang2022surveynested . Prior work has focused on general-domain NER devlin2019bert as well as clinical NER gerardin2022multilabel ; sun2021biomedical . This work can be mainly divided into two approaches, depending on the type of language models used.
The first approach is to use pre-trained Masked Language Models (MLM). This type of models is first pre-trained to predict randomly-selected masked words in large text corpora using a dense vector representation of every token (e.g., word) in the text devlin2019bert ; peters2018deep . Leveraging these models for NER usually involves training a linear projection to map vector representations into an NER tagging of the sentence, while jointly fine-tuning the parameters of the language model itself for the downstream task of NER. This approach has gained a lot of attention in the community since the release of the BERT architecture devlin2019bert , and became the go-to solution for building robust NER systems.
However, when dealing with clinical NER, this approach encounters two problems. First, due to the sensitive nature of EHRs, public corpora are rare, restrictively licensed, with limited availability in languages other than English. This leads the community to solutions built on MLMs pre-trained mainly on general-domain corpora, and thus suffering from the domain shift problem. Second, for fine-tuning to be efficient, large corpora of in-domain annotated text is required jia2019cross ; liu2021crossner . But clinical NER annotation campaigns are highly expensive and time-consuming due to the level of domain expertise needed to perform it luo2020n2c2 ; neveol2014quaero ; dogan2014ncbi ; baez2020chilean . Additionally, due to the diversity of clinical cases, data annotated for one biomedical application might not necessarily be helpful for another. Hence the need for data-efficient clinical NER, also known as few-shot NER.
The second, novel, approach is to use pre-trained Causal Language Models (CLM). These substantially larger models are pre-trained on (often larger) corpora as generative, auto-regressive models. That is, the model is given a series of tokens or prompt as input and estimates the most probable following series of tokens. To leverage these language models for downstream tasks such as NER, one can formulate the task in natural language in a prompt. The prompt is formulated in such a way that the continuity of the text involves resolving the task. The language model is then used to predict this continuity. This process is often called "In-context learning" (ICL) brown2020language . Optionally, one can build a prompt featuring a few demonstrations of the task resolved for other instances (in this case, task-specific NER-labeled instances), along with the new test input lee2022good . The model thus outputs an estimation of the most probable NER tagging for the test input instance. While MLMs have been studied for few-shot NER du2021representation , CLMs seem more naturally adapted to it. ICL learning has in fact proven particular success with LLMs in few-shot learning, showing state-of-the-art results in a wide set of NLP tasks shin2022effect ; wei2022chain ; srivastava2023beyond .
However, the superiority of CLMs over MLMs for NER is questionable. First, many efforts towards "few-shot learning" with CLMs design prompts based on their performance on large held-out validation datasets brown2020language ; tam2021improving ; radford2021learning ; qin2021learning . This is problematic because ICL is shown zhao2021calibrate ; lu2022fantastically ; min2022rethinking to depend greatly on the prompt structure : a small change in task phrasing, the examples presented, the order of examples, or the tagging format can affect the performance. Therefore, making these choices assuming large annotated validation dataset leads to results that are shown perez2021true to be over-optimistic and impossible to find in a real few-shot setting. Second, most of these studies have been mainly concentrated on the English language, and on GPT-based models wang2023gptner ; ashok2023promptner ; hu2023zero ; jimenez2022thinking . This can lead to over-fitting the prompts for this language and this language model. Hence, we identify a need for a systematic, model-independent, study of the prompt building in the clinical context, and for languages other than English. The contributions of this work are as follows:
  1. 1.
    We present and compare the most recent NER prompting techniques when applied to clinical NER in three languages : English, French and Spanish. To the best of our knowledge, this is the first work focused on NER prompts for languages other than English, and the first work comparing prompts for clinical NER.
  2. 2.
    We bring particular attention to tagging prompts, a novel NER prompting fashion, and measure the improvements brought by it.
  3. 3.
    We offer a fair comparison with the best-performing MLMs in a few-shot setting across languages, models, and prompt-structures when applied to few-shot clinical NER.
  4. 4.
    We conduct easily reproducible experiments, using easy-to-implement methods, exclusively on publicly available datasets and publicly available language models.
2 BACKGROUND AND SIGNIFICANCE
2.1 Few-shot NER with pre-trained Masked Language Models
The standard fashion of using MLMs for NER is as encoders. Usually, an NER tagging layer is trained from scratch to map the encoding of text into NER tagging of its tokens devlin2019bert . Other approaches have been proposed to leverage MLMs for few-shot learning. Namely, metric learning fritzler2019few ; yang2020simple ; huang2021few proposes to train systems to instead learn a metric over the output space. New instances can then be classified based on the distance separating them from other labeled instances. Label encoding aly2021leveraging ; ma2022label ; hou2020few suggests, instead, to leverage label names or textual label descriptions and encode them along with the instances in order to better tag them.
2.2 Few-shot NER with Causal Language Models
Recently, prompt construction has gained interest in the community brown2020language ; liu2023pre . While most related work focused on studying prompt formulation and exploring better-performing prompt structures wei2022chain ; ashok2023promptner ; vilar2023prompting ; wang2023gptner also known as "prompt engineering", other work proposed continuous optimization of the prompt through prompt tuning ma2022templatefree ; layegh2023contrastner ; hu2023vpn , usually reporting marginal improvements over baselines.
There is no standard, widely adopted manner of building NER prompts liu2023pre . In fact, NER associates to each instance a set of spans, each of which having a type. This structured nature of the prediction make it hard to find an intuitive but efficient manner to prompt a language model for NER, that adapts well to all contexts.
For instance, the main practice is to use separate prompts for different entity types li2020unified ; liu2022qaner ; chen2023learning . This choice seems well-suited when the task is interested in a handful of types of entities (typically 5-10). When interested in less entity types, a single prompt can be used for detecting all entities ashok2023promptner . On the other hand, if there is more entity types, it could be interesting to enumerate every possible span in the input sentence and let the model predict the entity type of the span, if any cui2021template . This method, on the inverse, is impractical for long inputs.
We identify three families of manners to prompt LLMs :
2.3 Few-shot clinical NER
MLM-based few-shot NER has also been explored in the biomedical domain geJBI2023 . Metric leaning yang2020simple and label encoding aly2021leveraging ; ma2022label have been explored, as well as other approaches such as active learning kormilitzin2021med7 , supervised pretraining huang2021baseline and MLM-based in-context learning lee2022good .
Few studies have focused on CLM-based few-shot clinical NER. In hu2023zero , GPT-3 and ChatGPT are evaluated on the 2010 i2b2/VA task uzuner20112010 in a zero-shot context. In jimenez2022thinking , GPT-3 is evaluated on a set of biomedical information extractions tasks including the NCBI-Disease dogan2014ncbi . Another interesting direction is partly fine-tuning liao2023parameter a general-domain CLM on clinical text han2023medalpaca ; toma2023clinical , and prompting the resulting CLM.
3 MATERIALS AND METHODS
3.1 Evaluation tasks
In order to evaluate the models. We use 14 publicly-available NER datasets, described as follows. For each study language, we selected two out-domain datasets and two or three in-domain datasets, aiming to use comparable resources (same genre, tagset, annotation guidelines) across languages whenever possible.
3.1.1 General-domain evaluation datasets
WikiNER nothman2013wikiner is a multilingual silver-standard annotated NER dataset. It consists in a late-2010 snapshot of Wikipedia in nine languages. Hyperlinks referring to persons, locations or organizations were automatically annotated. We use the English, French and Spanish versions of this dataset.
CoNLL-2002 sang2002conll and CoNLL-2003 sang2003conll are two manually-annotated multilingual NER dataset released as a part of CoNLL shared tasks. Mentions of persons, locations, organizations and miscellaneous entities are annotated. We use the Spanish data of the 2002 version, which is a collection of news wire articles made available by the Spanish EFE News Agency, released in May 2000. We use the English data of the 2003 version, which consists of Reuters news stories between 1996 and 1997.
Quaero French Press grouin-etal-2011-proposal is a manually annotated corpus of about 100 hours of speech transcribed from French speaking radio broadcast. This corpus was used in the 2011 Quaero named entity evaluation campaign. It comprises annotations for 5 entity types further divided into 32 subtypes. Our experiments relied on the five entity types: persons, locations, organizations, functions, and facilities.
3.1.2 Clinical evaluation datasets
E3C magnini2021e3c is a European multilingual corpus (Italian, English, French, Spanish, and Basque) of semantically annotated clinical narratives. The texts are collected from multiple publicly-available sources such as abstracts extracted from CC-licensed journals. We use the gold standard material available from the English, French and Spanish versions of this dataset. The clinical narratives are annotated with 6 entity types : actors, body parts, events, RMLs (measurements and test results) and clinical entities.
The n2c2-2019 luo2020n2c2 shared task focuses on medical concept normalization. It uses the MCN corpus developed by luo2019mcn , often referred to as the n2c2-2019 dataset. It includes discharge summaries from the Partners HealthCare and Beth Israel Deaconess Medical Center. In order to convert the medical concept normalization task into an NER task, we use the annotated Concept Unique Identifiers (CUIs) to map each mention to the corresponding UMLS semantic group lindberg1993umls ; mccray2001aggregating .
The NCBI-Disease dogan2014ncbi corpus gathers 793 PubMed abstracts where mentions of diseases are annotated in four types depending on their syntax : Specific Diseases (e.g. diastrophic dysplasia), Disease Classes (e.g. an autosomal recessive disease), Composite Mentions (e.g. colorectal, endometrial, and ovarian cancers), and Modifiers (e.g. C7-deficient) .
QuaeroFrenchMed neveol2014quaero consists of two text sources that we treat separately. The first part, EMEA is a collection of 13 patient information leaflets on marketed drugs from the European Medicines Agency (EMEA). The second part, MEDLINE, consist of 2,500 titles of research articles indexed in the MEDLINE database [^1] http://pubmed.ncbi.nlm.nih.gov/. The two parts are annotated with 10 entity types corresponding to UMLS semantic groups.
The Chilean Waiting List baez2020chilean corpus consists of 900 de-identified referrals for several specialty consultations in Spanish from the waiting list in Chilean public hospitals, manually annotated with 10 entity types : abbreviations, body parts, clinical findings, diagnostic procedure, diseases, family members, laboratory or test results, laboratory procedures, medications, procedures, signs or symptoms and therapeutic procedures. It can be noted that these types can be redundant (e.g. all diagnostic procedures are also annotated as procedures).
3.1.3 Few-shot learning set-up
In order to study our models in a few-shot context, we simulate the few-shot context by only providing the models with a few annotated examples. These are all the annotated examples models are allowed to use in training, in prompting and in validation. In this study, we choose to mainly focus on $k=100$ sentences, which corresponds to one to two hours of annotation in the clinical domain neveol2014quaero ; campillos2018french . We use a fixed random seed $p$ to choose $k$ examples among all those available in the actual corpus. In section 5.4.2, we discuss the effect of the choice of $k$ and the choice of $p$.
Additionally, we test the best-performing models with a full train dataset for a skyline comparison.
3.2 Language Models
Table LABEL:tab:LM_features presents an overview of the language models used in our study. While French and Spanish are covered in many of the causal models, we can observe that English is ubiquitous. Except for mBERT and XLM-RoBERTa, masked language models cover only one of our study languages.
3.3 NER with Masked Language Models
As mentioned in section 2.1, Masked Language Models have been adapted to few-shot learning in architectures suited for low-ressource contexts fritzler2019few ; yang2020simple ; huang2021few ; aly2021leveraging ; ma2022label ; hou2020few However, in this work, we are interested in comparing the novel CLM approaches to the widespread, standard MLMs usage in without any further adaptation for few-shot learning. We use NLStruct wajsburt2021these , an open-source Python library that implements the standard approach described in section 2.1. In addition, it processes nested entities, which are present in some or the study corpora. Instead of classifying the representation of every token separately into a BIO scheme, NLStruct classifies every representation span directly into entity types using a biLSTM-CRF tagger. We train the model for 20 epochs on 80% of the data and use the remaining held-out 20% for early stopping.
3.4 NER with Causal Language Models
Our experiments prompt models to tag entities in the input sentence, instead of listing them. We discuss this choice in further detail in section 5.4.1.
The upper part of figure 1 shows a sample tagging prompt, highlighting sections in the prompt that guided use to design features for prompt phrasing. The 9 prompt phrasing features we considered are described below.
Figure 1: Example of a tagging prompt, used in the main experiment (top) and a self-verification prompt (bottom) for detecting DISO mentions in n2c2-2019
  1. 1.
    Prompt language: By default, we prompt all language models in English, as it is the most ubiquitous language in all of their training corpora. This feature allows the model to be prompted in French or Spanish, to align the prompt language with that of the test sentence.
  2. 2.
    Additional sentences: By default, we present 5 annotated sentences in the prompts. This feature presents 5 additional sentences (i.e., 10 sentences in total).
  3. 3.
    Self verification: By default, we follow wang2023gptner by selecting the 5 closest sentences to the test sentence in terms of TF-IDF distance. The mentions tagged by the model are then considered to be the model’s final predictions. This feature selects instead the 5 sentences featuring the most entities of the target type and features them in an initial prompt. Intuitively, this prompt results in higher recall and lower precision. A second "self-verification" prompt is then used over the model’s initial predictions in order to filter out the false positives. A sample self-verification prompt is shown in the bottom part of figure 1. The number of demonstrations follows that of the main prompt.
  4. 4.
    Taggers: By default, we follow wang2023gptner prompting the model to surround mentions with @@ and ##. This feature prompts it to surround mentions with quotes << and >> instead.
  5. 5.
    Address a specialist in the prompt: By default, the first sentence is the task description shown in figure 1. This feature starts the prompt with You are an excellent <specialist>. You can identify all the mentions of <entity-type> in a sentence, by putting them in a specific format. Here are some examples you can handle: instead. The <specialist> is a linguist or a clinician, following the task domain.
  6. 6.
    Include label definitions in the prompt: This feature adds a one-sentence description for each entity type. Full entity descriptions used can be found in appendix 1.
  7. 7.
    Introductory sentence for the test instance: By default, the demonstrations are immediately followed by the test instance. This feature separates them with Identify all the mentions of <entity-type> in the following sentence, by putting <begin-tag> in front and a <end-tag> behind each of them.
  8. 8.
    Require a long answer for the self-verification: By default, the self-verification prompt demonstrates Yes (respectively No) as answers. This feature demonstrates <mention> is a(n) <entity-type>, yes. (respectively <mention> is not a(n) <entity-type>, no.) instead.
  9. 9.
    Dialogue template: This feature replaces the Input: and Output: in the prompt by dashes to imitate a dialog template.
ICL learning performance is shown to vary greatly depending on the exact phrasing of the prompt lu2022fantastically ; min2022rethinking . In addition, the optimal choice for each of these features can vary depending on the model used. For instance, intuitively, models that are heavily pretrained on the English language tend to perform better with an English template than one in the language of the corpus.
While our system aims to search for the best combination of parameters for each model, a grid search over them would require $5122^{9}=$ 512 experiments for each model, for each dataset. In order to build a lighter system, we choose to perform a greedy search. We iterate over the features in this order, testing the non-default value, and keeping it if it performs better than the default. In section 5.4.3, we compare this approach to a grid search for one model over one dataset.
Many efforts towards "few-shot learning" with CLMs optimize prompts on large held-out validation datasets brown2020language ; tam2021improving ; radford2021learning ; qin2021learning . This leads to results that are shown perez2021true to be over-optimistic. A fair comparison between MLMs and CLMs should compare them with access to the same (small) number of annotated instances, which corresponds to our $k=100$. In this no-training context, we follow perez2021true optimizing these features through a leave-one-out cross-validation (LOOCV).
3.5 Measures
We assess the performance of models using F-measure and grams of CO2 emissions.
  • 1.
    Micro-F1 for simplicity, we evaluate models over one global performance score. It is computed as the micro-average of F1-measures of the retrieval of each entity type.
  • 2.
    Carbon footprint we use GreenAlgorithms v2.2 lannelongue2021green  [^2] http://calculator.green-algorithms.org/ to estimate the carbon footprint of each experiment, based on factors such as runtime, computing hardware and location where electricity used by our computer facility was produced.
4 RESULTS
4.1 Performance
Table LABEL:tab:results and figure 2 describe the performance of the tested models.
4.2 Environmental Impact
Appendix 2 details the carbon emission estimations for all of our experiments. In particular, we estimate the experiment using Mistral-7B over ConLL-2003 to have generated 41g of CO2 equivalent. (6g for prompt optimization and 35g for inference on the test set). LLaMA-2-70B, around 10 times larger, is estimated to have generated 191g of CO2 equivalent. (44g for prompt optimization and 147g for inference on the test set).
On the other hand, the experiment on a the BERT-large MLM is estimated to have generated 6g of CO2 equivalent. (2g for fine-tuning and training and 4g for inference on the test set).
In total, the experiments described in this paper are estimated to have generated around 27kg of CO2 equivalent (25kg for the main experiments, and 2kg for ablation).
(a) English
(b) French
(c) Spanish
Figure 2: General vs. Clinical performance of studied models
5 DISCUSSION
5.1 Comparison of model performance
Our experiments offer a comparative analysis of various masked and causal language models for named entity recognition. We further focus the scope to low resource settings commonly found in real-life biomedical applications. Our results show that, despite being smaller and theoretically requiring a larger amount of training data, masked, "BERT-like" models consistently outperform CLMs in this context. In addition, this performance comes at a much lower environmental impact (CO2 emissions are 10-50 times lower for MLMs vs. CLMs).
Another important finding is that, in addition to their higher scores, the different MLMs achieve results that are relatively close to each other. For example, on the WikiNER generalist task in English, the 4 general-domain models tested achieved F1-scores of between 0.768 and 0.79.
Besides, we show that the MLMs specialized in the biomedical field (ClinicalBERT, CamemBERT-bio, etc.), on the one hand, suffer a sharp drop in general domain tasks, illustrating the classical issue of "catastrophic forgetting"; and on the other hand, do not bring any significant improvement in specialized tasks, with the exception of Spanish tasks. This comment must, however, be balanced by the difference in size between the models : all the specialized models only have 110 million parameters.
Named entity recognition based on BERT-type representations has received a great deal of attention in recent years, and is undoubtedly more mature than the use of CLMs for this task. We have implemented the CLM-based NER techniques recently published in the literature, to the best of our knowledge. It is, of course, possible that new approaches will make it possible to increase performance in the future. However, this is arguably a difficult task for a generative model, as it is highly constrained in its syntax and its evaluation. These results are no indication of performance on other tasks such as classification.
5.2 Practical use of language models for low-resource NER
Overall, our experiments suggest that the performance of language models for clinical named entity recognition is currently sub-optimal. In particular, even MLM-based models fail to approach the performance of fully supervised models. The three large models trained with the entirety of each training dataset (skylines Table LABEL:tab:results) systematically outperform the best few-shot results, by between 5% and 16% for the general domain, and between 8% and 48% for the biomedical domain. However, performance can be judged satisfactory enough for pre-annotation use, to complement or accelerate manual annotation, for example in an online or active learning context.
5.3 Limitations of our study
Random Noise
In MLM experiments, the parameters of the NER tagging layer added on top of the pretrained language model are initialized randomly. Likewise, in CLM experiments, the demonstrations in the prompts are shuffled randomly, and the negative examples in the self verification prompts are selected randomly. These random decisions can introduce noise in our performance measurements. Replicating all the experiments would allow us to draw more solid conclusions reimers-gurevych-EMNLP2017 , but would also come at a considerable cost (25kg of CO2 equivalent, and around 56 hours of computation for each replication). The large number of models tested and tasks addressed can however comfort the main observations of this article. For instance, we use Almost Stochastic Order (ASO)  [^3] Given the performance scores of two algorithms A and B, each of which run several times with different settings, ASO computes a test-specific value ($\epsilon_{\_}{min}$) that indicates how far algorithm A is from being significantly better than algorithm B. If distance $\epsilon_{\_}{min}$ = 0.0, one can claim that A stochastically dominant over B with the predefined significance level. The literature commonly interprets $\epsilon_{\_}{min}$ 0.5 as an indicator of a significant superiority of A over B. dror2019deep with a confidence level $\alpha=0.05$ to measure the significance of the superiority of MLMs over CLMs for each dataset separately. We do not always observe satisfying values of $\epsilon_{\_}{min}$ as to whether MLMs perform better than CLMs on general-domain NER (0.54, 0.121 and 0.019 respectively on WikiNER-English, CoNLL2003 and WikiNER-French). Regarding clinical NER, MLMs perform significantly better than CLMs : MLMs are stochastically dominant over CLMs ($\epsilon_{\_}{min}$=0) for all clinical datasets.
Data contamination.
The size of the training corpora used for creating LLMs makes it increasingly difficult to control for data contamination, i.e. the presence of test corpora. The community is calling for efforts towards better documentation of training datasets bender-friedman-2018-data . While some datasets are by construction incompatible with some models (e.g., there is no Spanish training corpus in GPT-J or LLAMA-2) we are unable to affirm full exclusion of all datasets from all models studied.
5.4 Ablation
To better understand the contribution of each step of our approach, we carried out a series of complementary experiments.
5.4.1 Listing prompts
In this section, we compare the adopted tagging prompts to listing prompts. In listing prompts, demonstrations simply list the tagged mentions. The list separator is optimized (in the same way as the taggers) between a comma and a newline character. Eventually, the introductory sentences asks to list entities. The results shown in table 1 further corroborate our choice of only focusing on tagging prompts.
English French Spanish Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL Listing prompts Mistral-7B 0.659 0.533 0.417 0.281 0.340 0.676 0.083 0.451 0.169 0.403 0.697 0.620 0.211 0.273 Tagging prompts Mistral-7B 0.754 0.646 0.488 0.291 0.395 0.727 0.428 0.590 0.229 0.333 0.720 0.707 0.083 0.374
Table 1: F1 scores obtained with the listing and tagging prompts.
5.4.2 Sample and sample size
We tested our approach with different samples and different sample sizes for one MLM : XLM-RoBERTa-large, and one CLM : Mistral-7B. The results are reported in table 2. It can be noted that, whereas the standard deviation with respect to $p$ is rather high, a significant difference can still be consistently observed between the two models across samples of the same size. We also observe that, as the number of annotated instances decreases, the performance of the MLM drops faster than that of the CLM.
CoNLL2003 n2c2
100 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.646 0.626 0.714 0.291 0.178 0.215
XLM-R-large 0.826 0.814 0.786 0.462 0.478 0.526
50 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.615 0.648 0.637 0.278 0.176 0.106
XLM-R-large 0.697 0.77 0.714 0.431 0.476 0.35
25 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.509 0.599 0.52 0.152 0.252 0.116
XLM-R-large 0.487 0.588 0.637 0.393 0.361 0.283
Table 2: F1 scores obtained over experiments with different training samples and different training sample sizes.
5.4.3 Hyperparameter grid search
In order to asses the quality of our adopted search method used to find the best feature combination to incorporate in the prompt, we compare this method to a naïve grid search over these features. We test all 512 combinations of our identified 9 features, for Mistral-7B over ConLL2003. The scores found through LOOCV vary between 0.0 and 0.656 with a mean value of 0.387 and a median of 0.46. The best-preforming combination is : Additional sentences, Self-verification, Introductory sentence for the test instance and Require a long answer for the self-verification, which is exactly the same combination we found initially through a greedy, tree search, that is around 20 times faster and less consuming.
6 CONCLUSION
This study assessed the performance of two types of large languages models, for few-shot entity recognition in three languages. Our experiments show that few-shot learning performance is significantly lower in the clinical vs. general domain. While masked language models perform better than causal language models (higher F1, lower CO2 emissions), few-shot use should be limited to assisting gold standard annotation rather than effective information extraction.
ACKNOWLEDGEMENT
This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-AD011014533). The authors thank Dr. Juan Manual Coria for his help phrasing prompts in Spanish.
Appendix A NER labels descriptions
Appendix B Carbon footprint