### v2

# Few-shot clinical entity recognition in English, French and Spanish: masked language models outperform generative model prompting

Abstract
Large language models (LLMs) have become the preferred solution for many natural language processing tasks. In low-resource environments such as specialized domains, their few-shot capabilities are expected to deliver high performance. Named Entity Recognition (NER) is a critical task in information extraction that is not covered in recent LLM benchmarks. There is a need for better understanding the performance of LLMs for NER in a variety of settings including languages other than English. This study aims to evaluate generative LLMs, employed through prompt engineering, for few-shot clinical NER. We compare 13 auto-regressive models using prompting and 16 masked models using fine-tuning on 14 NER datasets covering English, French and Spanish. While prompt-based auto-regressive models achieve competitive F1 for general NER, they are outperformed within the clinical domain by lighter biLSTM-CRF taggers based on masked models. Additionally, masked models exhibit lower environmental impact compared to auto-regressive models. Findings are consistent across the three languages studied, which suggests that LLM prompting is not yet suited for NER production in the clinical domain.
1 Introduction
Electronic Health Records (EHR) are rich sources of clinical information Demner-Fushman et al. (2009), which often appear in unstructured text only Escudié et al. (2017). Efficiently extracting information from EHRs into a more structured form can help advance clinical research, public health surveillance and clinical decision support Wang et al. (2018).
Named Entity Recognition (NER) is a critical primary step in information extraction that aims to identify and categorize mentions of relevant entities in text. In the context of clinical information extraction, these can be mentions of clinical entities such as disorders or drugs. Extracting these entities can significantly enhance concept normalization Cho et al. (2017); Wajsbürt et al. (2021); Sung et al. (2022) as well as facilitate interpreting patient profiling and phenotyping Gérardin et al. (2022). Clinical NER is widely regarded as a challenging problem : clinical entities are often jargon or ambiguous, and clinical texts have a nonstandard phrasal structure Luo et al. (2020); Leaman et al. (2015). Additionally, the sensitive nature of EHRs results in a lack of publicly available clinical corpora, which are often restrictively licensed and predominantly available in English. Moreover, the annotation of clinical NER data demands substantial domain expertise, rendering such campaigns both costly and time-intensive Luo et al. (2020); Névéol et al. (2014); Doğan et al. (2014); Báez et al. (2020). Additionally, due to the diversity of clinical cases, data annotated for one biomedical application might not necessarily be helpful for another. Consequently, there is a critical need for data-efficient clinical NER, specifically few-shot NER approaches.
Large Language Models (LLMs) - specifically, causal, generative models - have demonstrated significant promise in few-shot learning across a wide variety of tasks, including text classification, machine translation, and question answering Brown et al. (2020); Radford et al. (2019). This supports our main research question: Can few-shot learning performance of LLMs transfer to the task of named entity recognition?
This study aims to evaluate generative LLMs, employed through prompt engineering, for few-shot clinical NER from the perspective of F1 performance and environmental impact.
Challenges of few-shot NER
We identify methodological challenges with the evaluation of few-shot prompting for NER.
First, there is no standard, widely adopted manner of prompting LLMs for NER tasks Li et al. (2022); Shen et al. (2023); Wang et al. (2023b); Keraghel et al. (2024), resulting in significant challenges for reproducibility and variations in results that are difficult to interpret. Second, many efforts towards "few-shot learning" with LLMs design prompts based on their performance on large held-out validation datasets Brown et al. (2020); Tam et al. (2021); Radford et al. (2021); Qin and Eisner (2021), which is not consistent with a few-shot set-up. In addition, in-context learning performance is shown to depend greatly on the prompt structure: a small change in task phrasing, the examples presented, the order of examples, or the tagging format can affect the performance Zhao et al. (2021); Lu et al. (2022); Min et al. (2022). Therefore, making these choices assuming large annotated validation dataset leads to performances that are shown Perez et al. (2021) to be over-optimistic and impossible to find in a real few-shot setting. Third, Zaghir et al. (2024) show that the majority of recent studies employing prompt engineering in medical applications lack a non-prompt-related baseline, such as fine-tuned BERT-like Masked Language Models (MLMs), which complicates the accurate assessment of the performance of these LLMs. Finally, most of these studies are mainly concentrated on English, and based on GPT, which is mainly trained on English, Jimenez Gutierrez et al. (2022); Wang et al. (2023b); Ashok and Lipton (2023); Hu et al. (2023b); Zaghir et al. (2024), limiting the generalizability of evaluations to other languages. The contributions of this work are as follows:
  1. 1.
    We describe a systematic algorithm for creating and optimizing prompts for NER, bringing a particular attention to tagging prompts Wang et al. (2023b), a novel NER prompting technique that recently showed particular promise García-Barragán et al. (2024); Magron et al. (2024).
  2. 2.
    We evaluate our algorithm in a true few-shot setting, by allowing prompt optimization only on the few annotated instances through cross-validation. 14 NER tasks were evaluated, spanning 6 general-domain datasets and 8 clinical datasets, with a focus on English, French and Spanish.
  3. 3.
    We compare this approach, applied to 13 generative LLMs, to the standard fine-tuning approach, applied to 16 MLMs, both in terms of performance and environmental impact. We provide our code at github.com/marconaguib/autoregressive_ner
2 Few-shot & clinical NER
Few-shot NER with pre-trained MLMs
Leveraging MLMs for NER usually involves using them as encoders, and training a linear projection to map vector representations into an NER tagging of the sentence, while jointly fine-tuning the parameters of the language model itself for the downstream task of NER Devlin et al. (2019). This approach has been widely studied Liu (2019); Petroni et al. (2020); Joshi et al. (2020); Schweter and Akbik (2021). Wajsbürt (2021) propose a similar architecture, enhanced with an entity decoder, to iteratively predict entity spans in the input, allowing the model to detect nested entities.
Few-shot NER can be performed by simply training such systems with the limited data available. Other approaches have been proposed to leverage MLMs specifically in few-shot setting. Namely, metric learning Fritzler et al. (2019); Yang and Katiyar (2020); Huang et al. (2021a) proposes to train systems to instead learn a metric over the output space. New instances can then be classified based on the distance separating them from other labeled instances. Label encoding Aly et al. (2021); Ma et al. (2022a); Hou et al. (2020) suggests, instead, to leverage label names or textual label descriptions and encode them along with the instances in order to better tag them.
Few-shot NER with generative LLMs
Recently, prompt construction has gained interest in the community Brown et al. (2020); Liu et al. (2023). While most related work focused on studying prompt formulation and exploring better-performing prompt structures Wei et al. (2022); Ashok and Lipton (2023); Vilar et al. (2023); Wang et al. (2023b) also known as "prompt engineering", other work proposed continuous optimization of the prompt through prompt tuning Ma et al. (2022b); Layegh et al. (2023); Hu et al. (2023a), usually reporting marginal improvements over baselines.
There is no standard, widely adopted manner of building NER prompts Liu et al. (2023). In fact, NER associates to each instance a set of spans, each of which having a type. This structured nature of the prediction make it hard to find an intuitive but efficient manner to prompt a language model for NER, that adapts well to all contexts.
For instance, the main practice is to use separate prompts for different entity types Li et al. (2020); Liu et al. (2022); Chen et al. (2023a). This choice seems well-suited when the task is interested in a handful of types of entities (typically 5-10). When interested in less entity types, a single prompt can be used for detecting all entities Ashok and Lipton (2023). On the other hand, if there is more entity types, it could be interesting to enumerate every possible span in the input sentence and let the model predict the entity type of the span, if any Cui et al. (2021). This method, on the inverse, is impractical for long inputs.
We identify three strategies for prompting LLMs. Constrained prompting attempts to better formulate the NER task by constraining the generation to fill in specific hand-crafted templates, usually adapted to MLMs Cui et al. (2021); Shen et al. (2023); Ye et al. (2023); Schick and Schütze (2021). Listing prompts consist in simply making the language model predict the entities in a list Ashok and Lipton (2023). Tagging prompts were studied more recently by Wang et al. (2023b). They make the language model surround entity mentions with special tags.
Few-shot clinical NER
MLM-based few-shot NER has also been explored in the biomedical domain Ge et al. (2023). Metric leaning Yang and Katiyar (2020) and label encoding Aly et al. (2021); Ma et al. (2022a) have been explored, as well as other approaches such as active learning Kormilitzin et al. (2021), supervised pretraining Huang et al. (2021b) and prompt-based learning Lee et al. (2022); Chen et al. (2023b); Cui et al. (2024).
Few studies have focused on LLM-prompting-based few-shot clinical NER. In Hu et al. (2024), GPT-3 and ChatGPT are evaluated on the 2010 i2b2/VA task Uzuner et al. (2011) in a zero-shot context. In Jimenez Gutierrez et al. (2022), GPT-3 is evaluated on a set of biomedical information extractions tasks including the NCBI-Disease Doğan et al. (2014).
On languages other than English,Meoni et al. (2023) use InstructGPT-3 is used to build multilingual training corpora to train smaller models, and in Ateia and Kruschwitz (2023), ChatGPT is evaluated on an NER challenge focused on extracting medical procedures in Spanish.
Another interesting direction is partly fine-tuning Liao et al. (2023) a general-domain LLM on clinical text Han et al. (2023); Toma et al. (2023); Yang et al. (2024), and prompting the resulting LLM.
3 Named Entity Recognition Experiments
3.1 Evaluation tasks
We use 14 publicly-available NER datasets (described in the next section) to compare prompted causal models to fine-tuned masked language models in few shot settings. For each study language, we selected two out-domain datasets and two or three in-domain datasets, aiming to use comparable resources (same genre, tagset, annotation guidelines) across languages whenever possible. We use official training, validation, and test subsets when available; otherwise, we apply an 80%-10%-10% split for training, validation, and testing.
3.1.1 General-domain evaluation datasets
WikiNER Nothman et al. (2013) is a multilingual silver-standard annotated NER dataset. It consists of a late-2010 snapshot of Wikipedia in nine languages. Hyperlinks referring to persons, locations or organizations were automatically annotated. We use the English, French and Spanish versions of this dataset.
CoNLL-2002 Tjong Kim Sang (2002) and CoNLL-2003 Tjong Kim Sang and De Meulder (2003) are two manually-annotated multilingual NER dataset released as a part of CoNLL shared tasks. Mentions of persons, locations, organizations and miscellaneous entities are annotated. We use the Spanish data of the 2002 version, which is a collection of news wire articles made available by the Spanish EFE News Agency, released in May 2000. We use the English data of the 2003 version, which consists of Reuters news stories between 1996 and 1997.
Quaero French Press Grouin et al. (2011) is a manually annotated corpus of about 100 hours of speech transcribed from French speaking radio broadcast. This corpus was used in the 2011 Quaero named entity evaluation campaign. It comprises annotations for 5 entity types further divided into 32 subtypes. Our experiments relied on the five entity types: persons, locations, organizations, functions, and facilities.
3.1.2 Clinical evaluation datasets
E3C Magnini et al. (2021) is a European multilingual corpus (Italian, English, French, Spanish, and Basque) of semantically annotated clinical narratives. The texts are collected from multiple publicly-available sources such as abstracts extracted from CC-licensed journals. We use the gold standard material available from the English, French and Spanish versions of this dataset. The clinical narratives are annotated with 6 entity types : actors, body parts, events, RMLs (measurements and test results) and clinical entities.
The n2c2-2019 Luo et al. (2020) shared task focuses on medical concept normalization. It uses the MCN corpus developed by Luo et al. (2019), often referred to as the n2c2-2019 dataset. It includes de-identified discharge summaries from the Partners HealthCare and Beth Israel Deaconess Medical Center. In order to convert the medical concept normalization task into an NER task, we use the annotated Concept Unique Identifiers (CUIs) to map each mention to the corresponding UMLS semantic group Lindberg et al. (1993); McCray et al. (2001).
The NCBI-Disease Doğan et al. (2014) corpus gathers 793 PubMed abstracts where mentions of diseases are annotated in four types depending on their syntax : Specific Diseases (e.g. diastrophic dysplasia), Disease Classes (e.g. an autosomal recessive disease), Composite Mentions (e.g. colorectal, endometrial, and ovarian cancers), and Modifiers (e.g. C7-deficient) .
QuaeroFrenchMed Névéol et al. (2014) consists of two text sources that we treat separately. The first part, EMEA is a collection of 13 patient information leaflets on marketed drugs from the European Medicines Agency (EMEA). The second part, MEDLINE, consist of 2,500 titles of research articles indexed in the MEDLINE database [^1] http://pubmed.ncbi.nlm.nih.gov/. The two parts are annotated with 10 entity types corresponding to UMLS semantic groups.
The Chilean Waiting List Báez et al. (2020) corpus consists of 900 de-identified referrals for several specialty consultations in Spanish from the waiting list in Chilean public hospitals, manually annotated with 10 entity types : abbreviations, body parts, clinical findings, diagnostic procedure, diseases, family members, laboratory or test results, laboratory procedures, medications, procedures, signs or symptoms and therapeutic procedures. It can be noted that these types can be redundant (e.g. all diagnostic procedures are also annotated as procedures).
3.1.3 Few-shot learning set-up
We simulate the few-shot context by only providing the models with a few annotated examples that can be used in training, prompting and validation. No additional examples are made available. In this study, we choose to mainly focus on $k=100$ sentences, which corresponds to one to two hours of annotation in the clinical domain Névéol et al. (2014); Campillos et al. (2018). We use a fixed random seed $p$ to choose $k$ examples among all those available in the corpus. In Section A.1.1, we discuss the effect of the choice of $k$ and of $p$.
Additionally, we train the best-performing language models with the entire training dataset to provide a skyline comparison, i.e. performance of the models outside the few-shot setting.
3.2 Language Models
Table 5 (appendix A.2) presents an overview of the language models used in our study. English is covered in all causal language models, which is not the case for Spanish and French. Except for mBERT and XLM-RoBERTa, masked language models cover only one of our study languages.
3.3 NER with MLMs
Although MLMs have been adapted to few-shot learning in architectures suited for low-resource contexts (see section 2), we compare LLM prompting to the simple and standard MLM usage without any further adaptation for few-shot learning, as an accessible and easily-reproducible baseline.
We use NLStruct Wajsbürt (2021), an open-source Python library  [^2] https://github.com/percevalw/nlstruct that implements the standard fine-tuning approach. NLStruct uses the representations provided by the language model to encode the input, then employs a bidirectional LSTM decoder and a CRF to iteratively predict the entities present in the encoded input, as described by Gérardin et al. (2022). This approach allows NLStruct to effectively handle nested entities, which are prevalent in some of the study corpora.
We train the model for 20 epochs on 80% of the data and use the remaining held-out 20% for early stopping.
3.4 NER with generative LLM prompting
Our experiments prompt models to tag entities in the input sentence, instead of listing them. This choice is supported by further experiments reported in Section A.1.2. The upper part of Figure 1 shows a sample tagging prompt, highlighting sections in the prompt that guided the design of features for prompt phrasing.
Prompt features
We describe below the nine optional features that control the phrasing of the prompt, as well as the criteria for selecting the few-shot examples featured in the prompt.
Figure 1: Example of a tagging prompt, used in the main experiment (top) and a self-verification prompt (bottom) for detecting DISO mentions in n2c2-2019
  1. 1.
    Prompt language: By default, we prompt all language models in English, as it is the most ubiquitous language in all of their training corpora. This feature allows the model to be prompted in French or Spanish, to align the prompt language with that of the test sentence.
  2. 2.
    Additional sentences: By default, we present 5 annotated sentences in the prompts. This feature presents 5 additional sentences (i.e., 10 sentences in total). Section A.1.3 discusses adding more demonstations to the prompt.
  3. 3.
    Self verification: By default, we select the 5 closest sentences to the test sentence in terms of TF-IDF distance, among the training set. The mentions tagged by the model are then considered to be the model’s final predictions. This feature selects instead the 5 sentences featuring the most entities of the target type and features them in an initial prompt. Intuitively, this prompt results in higher recall and lower precision. A second "self-verification" prompt is then used over the model’s initial predictions in order to filter out the false positives. A sample self-verification prompt is shown in the bottom part of Figure 1.
    The number of demonstrations follows that of the main prompt.
  4. 4.
    Taggers: By default, we follow Wang et al. (2023b) prompting the model to surround mentions with @@ and ##. This feature prompts it to surround mentions with quotes << and >> instead.
  5. 5.
    Address a specialist in the prompt: By default, the first sentence is the task description shown in Figure 1. This feature starts the prompt with You are an excellent <specialist>. You can identify all the mentions of <entity-type> in a sentence, by putting them in a specific format. Here are some examples you can handle: instead. The <specialist> is a linguist or a clinician, following the task domain.
  6. 6.
    Include label definitions in the prompt: This feature adds a one-sentence description for each entity type. Full entity descriptions used can be found in appendix A.3.
  7. 7.
    Introductory sentence for the test instance: By default, the demonstrations are immediately followed by the test instance. This feature separates them with Identify all the mentions of <entity-type> in the following sentence, by putting <begin-tag> in front and a <end-tag> behind each of them.
  8. 8.
    Require a long answer for the self-verification: By default, the self-verification prompt demonstrates Yes (respectively No) as answers. This feature demonstrates <mention> is a(n) <entity-type>, yes. (respectively <mention> is not a(n) <entity-type>, no.) instead.
  9. 9.
    Dialogue template: This feature replaces the Input: and Output: in the prompt by dashes to imitate a dialog template.
Identification of optimal prompt configuration
In-context learning performance is shown to vary greatly depending on the exact phrasing of prompts Lu et al. (2022); Min et al. (2022). In addition, the optimal choice for each of these features can vary depending on the model used. For instance, intuitively, models that are heavily pretrained on the English language tend to perform better with an English template than one in the language of the corpus.
While our system aims to search for the best combination of parameters for each model, a grid search over them would require $5122^{9}=$ 512 experiments for each model, for each dataset. In order to build a lighter system, we opt for a greedy search. We iterate over the features and select options that perform better than the default. In Section A.1.4, we illustrate how the optimal parameter combination can be obtained with a greedy search using 5% of the computation required for grid search. Many efforts towards "few-shot learning" with LLMs design prompts on large held-out validation datasets Brown et al. (2020); Tam et al. (2021); Radford et al. (2021); Qin and Eisner (2021). This leads to results that are shown Perez et al. (2021) to be over-optimistic. A true few-shot evaluation involves optimization with access to a small number of annotated instances, which corresponds to our $k=100$. In this no-training context, we follow Perez et al. (2021) optimizing these features through a leave-one-out cross-validation (LOOCV) over the validation set.
3.5 Performance metrics
Micro-F1
For simplicity, we evaluate models over one global information extraction performance score. It is computed as the micro-average of F1-measures for each entity type.
Carbon footprint
We use GreenAlgorithms v2.2 Lannelongue et al. (2021)  [^3] http://calculator.green-algorithms.org/ to estimate the carbon footprint of each experiment, based on factors such as runtime, computing hardware and location where electricity used by our computer facility was produced.
4 Results
4.1 Environmental Impact
Figure 2: Carbon emission (g) incurred by resolving ConLL-2003 using three models
Figure 2 compares carbon emission incurred by resolving ConLL-2003 using LLM prompting (using Llama-2-70B and Mistral-7B) as well as with fine-tuning BERT-large. For comparison, the impact of using Llama-2-70B is comparable to that of driving a thermal car of average size for 1 kilometer.
Appendix A.4 details the carbon emission estimations for all of our experiments. In total, the experiments described in this paper are estimated to have generated around 31kg of CO2 equivalent (29kg for main experiments, and 2kg for ablation).
English French Spanish # Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL Few-shot approaches Causal 1 Llama-2-70B 0.728 0.721 0.312 0.309 0.400 0.740 0.400 0.483 0.201 0.312 0.805 0.616 0.021 0.339 2 Llama-3-8B-Instruct 0.756 0.727 0.478 0.252 0.390 0.796 0.508 0.652 0.443 0.433 0.784 0.764 0.226 0.437 3 Mistral-7B 0.754 0.646 0.488 0.291 0.395 0.727 0.428 0.590 0.229 0.333 0.720 0.707 0.083 0.374 4 Phi-3-medium-instruct 0.464 0.302 0.234 0.112 0.178 0.394 0.322 0.317 0.139 0.161 0.612 0.461 0.045 0.280 5 BLOOM-7B1 0.524 0.557 0.279 0.113 0.151 0.148 0.206 0.320 0.197 0.120 0.470 0.419 0.051 0.117 6 Falcon-40B 0.686 0.708 0.280 0.279 0.305 0.662 0.456 0.378 0.279 0.283 0.720 0.543 0.072 0.267 7 GPT-J-6B 0.521 0.493 0.167 0.179 0.238 0.423 0.244 0.334 0.080 0.177 0.005 0.142 0.021 0.162 8 OPT-66B 0.608 0.495 0.227 0.157 0.234 0.624 0.406 0.019 0.206 0.283 0.166 0.273 0.043 0.204 9 Vicuna-13B 0.657 0.708 0.355 0.236 0.300 0.677 0.350 0.399 0.207 0.326 0.744 0.250 0.040 0.213 10 Vicuna-7B 0.594 0.489 0.259 0.147 0.172 0.591 0.277 0.439 0.152 0.296 0.659 0.569 0.042 0.151 11 BioMistral-7B 0.414 0.354 0.175 0.086 0.257 0.547 0.299 0.350 0.236 0.186 0.578 0.540 0.014 0.226 12 Medalpaca-7B 0.537 0.586 0.272 0.138 0.132 0.529 0.142 0.259 0.162 0.252 0.581 0.490 0.088 0.220 13 Vigogne-13B 0.593 0.655 0.252 0.176 0.309 0.515 0.250 0.464 0.099 0.142 0.580 0.561 0.010 0.198 Masked 14 mBERT 0.768 0.804 0.624 0.378 0.401 0.801 0.728 0.741 0.588 0.428 0.812 0.760 0.324 0.432 15 XLM-R-large 0.786 0.826 0.637 0.462 0.471 0.811 0.781 0.762 0.629 0.531 0.797 0.781 0.325 0.528 16 BERT-large 0.776 0.814 0.626 0.435 0.422 - - - - - - - - - 17 RoBERTa-large 0.790 0.829 0.626 0.462 0.552 - - - - - - - - - 18 Bio_ClinicalBERT 0.528 0.542 0.621 0.469 0.420 - - - - - - - - - 19 ClinicalBERT 0.462 0.597 0.622 0.480 0.397 - - - - - - - - - 20 MedBERT 0.613 0.673 0.607 0.478 0.504 - - - - - - - - - 21 CamemBERT-large - - - - - 0.829 0.793 0.768 0.661 0.564 - - - - 22 FlauBERT-large - - - - - 0.826 0.778 0.760 0.635 0.540 - - - - 23 DrBERT-4GB - - - - - 0.587 0.599 0.730 0.602 0.497 - - - - 24 CamemBERT-bio - - - - - 0.782 0.761 0.779 0.636 0.557 - - - - 25 BETO - - - - - - - - - - 0.794 0.732 0.352 0.522 26 PatanaBERT - - - - - - - - - - 0.802 0.769 0.343 0.487 27 TulioBERT - - - - - - - - - - 0.804 0.798 0.340 0.482 28 BSC-BioEHR - - - - - - - - - - 0.804 0.758 0.354 0.578 29 BSC-Bio - - - - - - - - - - 0.804 0.775 0.358 0.552 Masked fully-supervised (skyline) RoBERTa-large 0.919 0.939 0.718 0.712 0.815 - - - - - - - - - CamemBERT-large - - - - - 0.928 0.834 0.828 0.748 0.713 - - - - BETO - - - - - - - - - - 0.918 0.881 0.411 0.736
Table 1: This table presents the micro-F1 obtained from few-shot experiments. Skyline results are obtained using all training data available instead of the few-shot setting.
4.2 Comparison of model performance
Table 1 presents the micro-F1 performance of the models on the test set of each dataset. Figures 3 to 5 present the results for each language with a vizualization of model type (circles for causal models vs. squares for masked models), model size in terms of parameters and domain (clinical in green vs. general in blue). Overall, results show that, despite being smaller and theoretically requiring a larger amount of training data, masked, "BERT-like" models consistently outperform generative LLM prompting in the context of few-shot NER commonly found in biomedical applications. Specifically, on English and Spanish, Figures 3 and 5, show that, while some generative LLMs are competitive in the general domain, they suffer a sharper performance drop in clinical NER, compared to fine-tuned MLMs. On French, Figure 4 suggests that generative LLM prompting does not seem competitive on either domain.
Moreover, this performance comes at a much lower environmental impact (CO2 emissions are 10-50 times lower for MLMs vs. prompted LLMs). Another important finding is that, in addition to high performance scores, MLMs achieve results that are relatively close to each other. For example, on the WikiNER generalist task in English, the 4 general-domain models tested achieved F1-scores of between 0.768 and 0.79.
Figure 3: Performance of models on English. The general performance is the average of micro-F1 obtained on WikiNER-en and CoNLL-2003. The clinical performance is the average on E3C-en, n2c2 and NCBI-Disease. The red lines represent the skyline performance obtained with the entirety of each training dataset.
Figure 4: Performance of models on French. The general performance is the average of micro-F1 obtained on WikiNER-fr and QuaeroFrenchPress. The clinical performance is the average on E3C-fr, EMEA and MEDLINE. The red lines represent the skyline performance obtained with the entirety of each training dataset.
Figure 5: Performance of models on Spanish. The general performance is the average of micro-F1 obtained on WikiNER-es and CoNLL-2002. The clinical performance is the average on E3C-es and CWL. The red lines represent the skyline performance obtained with the entirety of each training dataset.
Besides, we show that domain-adapted MLMs (e.g., ClinicalBERT, CamemBERT-bio) exhibit a sharp performance drop in general domain tasks, illustrating the classical issue of "catastrophic forgetting". In addition, they do not contribute performance improvements in specialized tasks, with the exception of Spanish tasks. However, there is a notable difference in model size: specialized models only have 110 million parameters (vs. 340 million for other models).
Named entity recognition based on BERT-type representations has received a great deal of attention in recent years, and is undoubtedly more mature than the use of LLM prompting for this task. We have implemented the prompt-based NER techniques recently published in the literature, to the best of our knowledge. It is, of course, possible that new approaches will make it possible to increase performance in the future. However, this is arguably a difficult task for a generative model, as it is highly constrained in its syntax and its evaluation. These results are no indication of performance on other tasks such as classification.
4.3 Practical use of language models for low-resource NER
Overall, our experiments suggest that the performance of language models for clinical named entity recognition is currently sub-optimal. In particular, even MLM-based models, simply fine-tuned on the limited data available, fail to approach the performance of fully supervised models. The three large models trained with the entirety of each training dataset (skylines Table 1) systematically outperform the best few-shot results, by 5% to 16% for the general domain, and 8% to 48% for the biomedical domain. However, performance can be judged satisfactory enough for pre-annotation use, to complement or accelerate manual annotation, for example in an online or active learning context.
5 Conclusion
This study assessed the performance of two types of large languages models, for few-shot entity recognition in three languages. Our experiments show that few-shot learning performance is significantly lower in the clinical vs. general domain. While masked language models perform better than causal language models (higher F1, lower CO2 emissions), LLM prompting is not yet suited for effective information extraction.
Limitations of our study
Random Noise and Significance
In MLM experiments, the parameters of the NER tagging layer added on top of the pretrained language model are initialized randomly. Likewise, in LLM prompting experiments, the demonstrations in the prompts are shuffled randomly, and the negative examples in the self verification prompts are selected randomly. These random decisions can introduce noise in our performance measurements. Replicating all the experiments would allow us to draw more solid conclusions Reimers and Gurevych (2017), but would also come at a considerable cost (29kg of CO2 equivalent, and around 64 hours of computation for each replication). The large number of models tested and tasks addressed can however support the main observations of this article. For instance, we use Almost Stochastic Order (ASO)  [^4] Given the performance scores of two algorithms A and B, each of which run several times with different settings, ASO computes a test-specific value ($\epsilon_{min}$) that indicates how far algorithm A is from being significantly better than algorithm B. If distance $\epsilon_{min}$ = 0.0, one can claim that A stochastically dominant over B with the predefined significance level. The literature commonly interprets $\epsilon_{min}$ 0.5 as an indicator of a significant superiority of A over B. Dror et al. (2019) with a confidence level $\alpha=0.05$ to measure the significance of the superiority of fine-tuned MLMs over prompted LLMs for each dataset separately. We do not always observe satisfying values of $\epsilon_{min}$ as to whether MLMs perform better than prompted LLMs on general-domain NER (0.426, 0.081 and 0.028 respectively on WikiNER-English, CoNLL2003 and WikiNER-French). Regarding clinical NER, MLMs perform significantly better than prompted LLMs : MLMs are stochastically dominant over prompted LLMs ($\epsilon_{min}$=0) for all clinical datasets.
Data contamination
The size of the training corpora used for creating LLMs makes it increasingly difficult to control for data contamination, i.e. the presence of test corpora. Moreover, Balloccu et al. (2024) describe the problem of "indirect contamination", encountered when models are iteratively improved by using data coming from users, including test dataset.
The community is calling for efforts towards better documentation of training datasets Bender and Friedman (2018). While some datasets are by construction incompatible with some models (e.g., there is no Spanish training corpus in GPT-J or Llama-2) we are unable to affirm full exclusion of all datasets from all models studied.
Fine-tuning generative models
Our study focuses on low-data scenarios prevalent in real-world biomedical applications, where computational resources, notably GPU availability, are constrained. Additionally, the use of models via API calls is challenging due to privacy and security concerns. Consequently, we selected methods requiring limited computing power : either fine-tuning "small", BERT-like models, or prompting larger, frozen, models. This pragmatic choice stems from the need to balance performance with computational efficiency, ensuring that the proposed methods remain feasible for implementation in resource-limited settings.
However, more computationally expensive solutions may exist and may be effectively deployed in other environments. Typically, one could use the limited available data to partially fine-tune a generative LLM for NER Liao et al. (2023).
Error analysis
To facilitate comparison between models, we evaluated models over one global information extraction performance score: the micro-F1 measure. Overall, precision and recall scores (not shown) are balanced across entity types. In clinical datasets, some entity types, such as chemicals and anatomy, yield higher performance than others such as devices and procedures. This disparity cannot be explained by the distribution in the training corpus, since prompts provide the same number of examples across entity types. We hypothesize that the regularity and prevalence in LLM pretaining corpora of some entities may explain the higher performance. In future work, we plan more in-depth analysis, to better understand the performance of the LLM prompting approach.
Ethical considerations
All the datasets analyzed in our experiments are publicly available corpora, used consistently with the relevant data use agreements.
The use of LLMs can incur significant environmental impact. We have measured carbon emissions using GreenAlgorithms. Our experiments support the conclusion that MLMs yield higher information extraction performance with lower carbon emission, compared to LLM prompting for NER.
Acknowledgements
This work has received funding from the French "Agence Nationale pour la Recherche" under grant agreement CODEINE ANR-20-CE23-0026-01. This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-AD011014533). The authors thank Dr. Juan Manual Coria for his help phrasing prompts in Spanish.
Appendix A Appendix
A.1 Ablation
To better understand the contribution of each step of our approach, we carried out a series of complementary experiments.
A.1.1 Sample and sample size
We tested our approach with different samples and different sample sizes for one MLM : XLM-RoBERTa-large, and one prompted LLM : Mistral-7B. The results are reported in table 2. It can be noted that, whereas the standard deviation with respect to $p$ is rather high, a significant difference can still be consistently observed between the two models across samples of the same size. We also observe that, as the number of annotated instances decreases, the performance of the MLM drops faster than that of the prompted LLM.
CoNLL2003 n2c2
100 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.646 0.626 0.714 0.291 0.178 0.215
XLM-R-large 0.826 0.814 0.786 0.462 0.478 0.526
50 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.615 0.648 0.637 0.278 0.176 0.106
XLM-R-large 0.697 0.77 0.714 0.431 0.476 0.35
25 annotated instances
p=1 p=2 p=3 p=1 p=2 p=3
Mistral-7B 0.509 0.599 0.52 0.152 0.252 0.116
XLM-R-large 0.487 0.588 0.637 0.393 0.361 0.283
Table 2: F1 scores obtained over experiments with different training samples and different training sample sizes.
A.1.2 Listing prompts
In this section, we compare the adopted tagging prompts to listing prompts. In listing prompts, demonstrations simply list the tagged mentions. The list separator is optimized (in the same way as the taggers) between a comma and a newline character. Eventually, the introductory sentences asks to list entities. The results shown in table 3 further corroborate our choice of only focusing on tagging prompts.
English French Spanish Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL Listing prompts Mistral-7B 0.659 0.533 0.417 0.281 0.340 0.676 0.083 0.451 0.169 0.403 0.697 0.620 0.211 0.273 Tagging prompts Mistral-7B 0.754 0.646 0.488 0.291 0.395 0.727 0.428 0.590 0.229 0.333 0.720 0.707 0.083 0.374
Table 3: F1 scores obtained with the listing and tagging prompts.
A.1.3 Number of demonstrations
The decision to limit the number of annotated examples presented to generative models in the prompt to a maximum of 10 is dictated by two constraints.
Firstly, models impose a limit on the number of input tokens over which attention is calculated. Most models used have a limit of 2048 tokens, but more permissive models, such as Mistral-7B, allow up to 8096 tokens. This constraint translates into a limit on the number of sentences that can be presented in the prompt, ranging from 40 to 50 for Mistral-7B and from 10 to 15 for less permissive models, depending on the corpora and tokenizers. For instance, consider the task of detecting body parts in the French section of E3C. Using Mistral-7B (and its tokenizer), the practical limit is around 40 examples, resulting in an average prompt of 7779.5 tokens. Using Vicuna-13B (and its tokenizer), the limit is around 11 examples, resulting in an average of 1851.5 tokens in the prompt.
Secondly, the improvement brought by adding more examples does not appear to be significant, as observed in Table 4, which shows the results obtained with Mistral-7B when the number of annotated examples is tripled. It is noteworthy that the marginal improvements achieved by tripling the number of examples come at a considerable cost, especially given the quadratic complexity with respect to the length of the prompt.
Therefore, we choose to limit the prompt to 5-10 examples, selected based on the chosen criteria (TF-IDF proximity to the reference sentence or the number of entities present). Instead of selecting examples independently, Gupta et al. (2023) propose selecting them interdependently to improve the representativeness of the prompt. This approach would be interesting to implement in our system in future work.
English French Spanish Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL 5/10 demonstrations Mistral-7B 0.754 0.646 0.488 0.291 0.395 0.727 0.428 0.590 0.229 0.333 0.720 0.707 0.083 0.374 15/30 demonstrations Mistral-7B 0.763 0.692 0.453 0.263 0.377 0.782 0.355 0.587 0.237 0.396 0.785 0.751 0.163 0.413
Table 4: F1 scores obtained with 5/10 demonstations vs. with 15/30 demonstations
A.1.4 Hyperparameter grid search
In order to assess the quality of our adopted search method used to find the best feature combination to incorporate in the prompt, we compare this method to a naïve grid search over these features. We test all 512 combinations of our identified 9 features, for Mistral-7B over ConLL2003. The scores found through LOOCV vary between 0.0 and 0.656 with a mean value of 0.387 and a median of 0.46. The best-preforming combination is : Additional sentences, Self-verification, Introductory sentence for the test instance and Require a long answer for the self-verification, which is exactly the same combination we found initially through a greedy, tree search, that is around 20 times faster and less consuming.
A.2 Evaluated models
Table 5 specifies relevant information about the tested models. In the training corpus column, we point out the data contaminations we know of. For instance, MedBERT’s documentation explicitly mentions N2C2 as part of its training data. This might lead to artificially high evaluation metrics, as the model is not generalizing to unseen data, but instead leveraging pre-learned information, which compromises the validity of the results in real-world applications.
# Model Number of parameters Training data size Training corpus Causal 1 Llama-2-70B[en] Touvron et al. (2023) 70B 2 trillion tokens A mix of publicly available online data, mainly in English 2 Llama-3-8B-Instruct[en] 8B over 15 trillion tokens "A new mix" of publicly available online data, mainly in English 3 Mistral-7B[?] Jiang et al. (2023) 7B Undisclosed Undisclosed 4 Phi-3-medium-instruct[en] Abdin et al. (2024) 14B 4.8 trillion tokens A combination of publicly available corpora, synthetic data and chat format supervised data, mainly in English 5 BLOOM-7B1[en][fr][es] Workshop et al. (2022) 7B 1.6 TB ROOTS Laurençon et al. (2022), a mix of datasets and pseudo-crawled data 59 languages 6 Falcon-40B[en][fr][es] 40B 1 trillion tokens RefinedWeb Penedo et al. (2023), a dataset of filtered and deduplicated web data 7 GPT-J-6B[en] Wang and Komatsuzaki (2021) 6B 825 GiB The Pile Gao et al. (2020), a mixture of public datasets and web data in English 8 OPT-66B[en] Zhang et al. (2022) 66B 180 billion tokens Crawled data from the web, mainly in English 9 Vicuna-13B[en]* Zheng et al. (2023) 13B 125K conversations Llama 2, fine-tuned on conversations collected from ShareGPT.com, mainly in English 10 Vicuna-7B[en]* Zheng et al. (2023) 7B 125K conversations Llama 2, fine-tuned on conversations collected from ShareGPT.com, mainly in English 11 BioMistral-7B[en]* Labrak et al. (2024) 7B 3 billion tokens Mistral, fine-tuned on the PMC Open Access Subset 12 Medalpaca-7B[en]* Han et al. (2023) 7B 400K Q.A. pairs Llama 2, fine-tuned on semi-generated medical question-answer pairs in English 13 Vigogne-13B[fr][en]* 13B 52K instructions Llama 2, fine-tuned on English instructions automatically translated to French Masked 14 mBERT[en][fr][es] Devlin et al. (2019) 110M Undisclosed A corpus featuring 104 languages built from undisclosed sources 15 XLM-R-large[en][fr][es] Conneau et al. (2020) 355M 2.5 TB Filtered CommonCrawl data containing 100 languages 16 BERT-large[en] Devlin et al. (2019) 345M 3,3 billion words BookCorpus Zhu et al. (2015), a dataset consisting of unpublished books and English Wikipedia. 17 RoBERTa-large[en] Liu et al. (2019) 355M 160 GiB BooksCorpus Zhu et al. (2015), English Wikipedia, and crawled web data 18 Bio_ClinicalBERT[en] Alsentzer et al. (2019) 110M 2 million clinical notes MIMIC-III Johnson et al. (2016), a database containing electronic health records from hospitalized ICU patients 19 ClinicalBERT[en] Wang et al. (2023a) 110M 1.2 billion words A large multi-center dataset with a corpus built from undisclosed sources 20 MedBERT[en] Vasantharajan et al. (2022) 110M 57 million words Community datasets (including N2C2 Luo et al. (2020)) and Crawled medical-related articles from Wikipedia 21 CamemBERT-large[fr] Martin et al. (2020) 335M 64 billion tokens OSCAR Suárez et al. (2020), a corpus of web data in French 22 FlauBERT-large[fr] Le et al. (2020) 335M 13 billion tokens A mix of French Wikipedia, French books, and French web data 23 DrBERT-4GB[fr] Labrak et al. (2023) 110M 1 billion words A mix of publicly available biomedical corpora in French (including QuaeroFrenchMed Névéol et al. (2014)). 24 CamemBERT-bio[fr] Touchent and de la Clergerie (2024) 110M 413 million words A mix of publicly available biomedical corpora in French (including E3C Magnini et al. (2021)). 25 BETO[es] Cañete et al. (2020) 110M 3 billion words Spanish Wikipedia and Spanish data from OPUS Tiedemann (2012) 26 PatanaBERT[es] 110M Undisclosed Spanish 27 TulioBERT[es] 110M Undisclosed Spanish 28 BSC-BioEHR[es] Carrino et al. (2022) 110M 1.1 billion tokens A mixture of biomedical community datasets including EHR documents and crawled data in Spanish 29 BSC-Bio[es] Carrino et al. (2022) 110M 963 million tokens A mixture of biomedical community datasets and crawled data in Spanish
Table 5: Characterization of the language models used in our experiments in terms of parameters and training corpus. Models marked with [en] (respectively [fr], [es]) are heavily trained on English (respectively French, Spanish). CLMs marked with * are fine-tuned versions of other CLMs.
A.3 NER labels descriptions
Tables 6 to 10 definitions of all the labels present in the studied datasets. These definitions were drawn from available annotation schemes and further curated by the authors and native speakers. These definitions were included in the prompt when the "Include label definitions in the prompt" feature was activated.
Tag Tag name (in singular) Description PER person names (a person’s name) These are names of persons such as real people or fictional characters. FAC facilities (a facility) These are names of man-made structures such as infrastructure, buildings and monuments. LOC locations (a location) These are names of geographical locations such as landmarks, cities, countries and regions. ORG organizations (an organization) These are names of organizations such as companies, agencies and political parties. FUNC functions and jobs (a function or a job) These are words that refer to a profession or a job. ACTI activities and behaviors (an activity or behavior) These are words that refer to human activities, behaviors or events as well as governmental or regulatory activities. ANAT anatomy (an anatomy) These are words that refer to the structure of the human body, its organs and their position, such as body parts or organs, systems, tissues, cells, body substances and embryonic structures. CHEM chemicals and drugs (a chemical or a drug) These are words that refer to a substance or composition that has a chemical characteristic, especially a curative or preventive property with regard to human or animal diseases, such as drugs, antibiotics, proteins, hormones, enzymes and hazardous or poisonous substances. CONC concepts and ideas (a concept or an idea) These are words that refer to a concept or an idea, such as a classification, an intellectual product, a language, a law or a regulation. DEVI medical devices (a device) These are words that refer to a medical device used to administer care or perform medical research. DISO disorders (a disorder) These are words that refer to an alteration of morphology, function or health of a living organism, animal or plant, such as congenital abnormalities, dysfunction, injuries, signs or symptoms or observations. GENE genes and molecular sequences (a gene or a molecular sequence) These are words that refer to a gene, a genome or a molecular sequence. GEOG geographical areas (a geographical area) These are words that refer to a country, a region or a city. LIVB living beings (a living being) These are words that refer to a living being or a group of living beings, such as a person or a group of persons, a plant or a category of plants, an animal or a category of animals. OBJC objects (an object) These are words that refer to anything animate or inanimate that affects the senses, such as physical manufactured objects. OCCU occupations (an occupation) These are words that refer to a professional occupation or discipline. ORGA organizations (an organization) These are words that refer to an organization such as healthcare related organizations. PHEN phenomema (a phenomemon) These are words that refer to a phenomenon that occurs naturally or as a result of an activity, such as a biologic function. PHYS physiology (a physiology) These are words that refer to any element that contributes to the mechanical, physical and biochemical functioning or organization of living organisms and their components. PROC procedures (a procedure) These are words that refer to an activity or a procedure that contributes to the diagnosis or treatment of patients, the information of patients, the training of medical personnel or biomedical research. EVENT events (an event) These are words that refer to actions, states, and circumstances that are relevant to the clinical history of a patient such as pathologies and symptoms, or more generally words like "enters", "reports" or "continue".
Table 6: Description of the NER tags used in our experiments for English.
Tag Tag name (in singular) Description TIMEX3 time expressions (a time expression) These are time expressions such as dates, times, durations, frequencies, or intervals. RML results and measurements (a result or a measurement) These are test results, results of laboratory analyses, formulaic measurements, and measure values. ACTOR actors (an actor) These are words that refer patients, healthcare professionals, or other actors relevant to the clinical history of a patient. Abbreviation abbreviations (an abbreviation) These are words that refer to abbreviations. Body_Part body parts (a body part) These are words that refer to organs and anatomical parts of persons. Clinical_Finding clinical findings (a clinical finding) These are words that refer to observations, judgments or evaluations made about patients. Diagnostic_Procedure diagnostic procedures (a diagnostic procedure) These are words that refer to tests that allow determining the condition of the individual. Disease diseases (a disease) These are words that describe an alteration of the physiological state in one or several parts of the body, due to generally known causes, manifested by characteristic symptoms and signs, and whose evolution is more or less predictable. Family_Member family members (a family member) These are words that refer to family members. Laboratory_or _Test_Result laboratory or test results (a laboratory or test result) These are words that refer to any measurement or evaluation obtained from a diagnostic support examination. Laboratory_Procedure laboratory procedures (a laboratory procedure) These are words that refer to tests that are performed on various patient samples that allow diagnosing diseases by detecting biomarkers and other parameters. Blood, urine, and other fluids and tissues that use biochemical, microbiological and/or cytological methods are considered. Medication medications (a medication) These are words that refer to medications or drugs used in the treatment and/or prevention of diseases, including brand names and generics, as well as names for groups of medications. Procedure procedures (a procedure) These are words that refer to activities derived from the care and care of patients. Sign_or_Symptom signs or symptoms (a sign or symptom) These are words that refer to manifestations of a disease, determined by medical examination or perceived and expressed by the patient. Therapeutic_Procedure therapeutic procedures (a therapeutic procedure) These are words that refer to activities or treatments that are used to prevent, repair, eliminate or cure the individual’s disease. CompositeMention composite mentions of diseases (a composite mention of diseases) These are words that refer to mentions of multiple diseases, such as "colorectal, endometrial, and ovarian cancers". DiseaseClass disease classes (a disease class) These are words that refer to classes of diseases, such as "an autosomal recessive disease". Modifier modifiers (a modifier of diseases) These are words that refer to modifiers of diseases, such as "primary" or "C7-deficient". SpecificDisease diseases (a disease) These are words that refer to specific diseases, such as "diastrophic dysplasia".
Table 7: Description of the NER tags used in our experiments for English, continued.
Tag Tag name (in singular) Description PER de noms de personnes (un nom de personne) Il s’agit des noms de personnes, qu’elles soient réelles ou fictives. FAC de productions humaines (une production humaine) Il s’agit des noms de structures faites par les humains comme des infrastructures, des bâtiments ou des monuments. LOC de lieux (un lieu) Il s’agit des noms de lieux comme des endroits, villes, pays ou régions. ORG d’organisations (une organisation) Il s’agit des noms d’organisations comme des entreprises, des agences ou des partis politiques. FUNC de fonctions et métiers (une fonction ou un métier) Il s’agit de mots qui se rapportent à une activité professionnelle. ANAT d’anatomie (une partie du corps) Il s’agit d’une entité se rapportant à la structure du corps humain, ses organes et leur position. Il s’agit principalement des parties du corpus ou organes, des appareils, des tissus, des cellules, des substances corporelles et des organismes embryonaires. CHEM de médicaments et substances chimiques (un médicament ou une substance chimique) Il s’agit d’une substance ou composition présentant des propriétés chimiques caractéristiques, en particulier des propriétés curatives ou préventives à l’égard des maladies humaines ou animales. Il s’agit principalement des médicaments disponibles en pharmacie, des antibiotiques, des proteines, des hormones, des substances dangereuses, des enzymes. DEVI de matériel (un matériel) Il s’agit d’un matériel utilisé pour administrer des soins ou effectuer des recherches médicales. DISO de problèmes médicaux (un problème médical) Il s’agit d’une altération de la morphologie, des fonctions, ou de la santé d’un organisme vivant, animal ou végétal. Il peut s’agir de malformations, de maladies, de blessure, de signe ou symptome ou d’une observation. GEOG de zones géographiques (une zone géographique) Il s’agit d’un pays, une région, ou une ville. LIVB d’êtres vivants (un être vivant) Il s’agit d’un être vivant ou groupe d’êtres vivants. Il peut s’agir d’une personne ou d’un groupe de personnes, d’une plante ou d’une catégorie de végétaux, d’un animal ou d’une catégorie d’animaux. OBJC d’objets (un objet) Il s’agit de tout ce qui, animé ou inanimé, affecte les sens. Ici, il s’agit principalement d’objets physiques manufacturés. PHEN de phénomènes (un phénomène) Il s’agit d’un phénomène qui se produit naturellement ou à la suite d’une activité. Il s’agit principalement de fonctions biologiques. PHYS de physiologie (une physiologie) Il s’agit de tout élément contribuant au fonctionnement ou à l’organisation mécanique, physique et biochimique des organismes vivants et de leurs composants. PROC de procédures (une procédure) Il s’agit d’une activité ou procédure contribuant au diagnostic ou au traitement des patients, à l’information des patients, la formation du personnel médical ou à la recherche biomédicale. EVENT d’événements (un événement) Il s’agit d’une action, d’un état ou d’une circonstance qui est pertinent pour l’histoire clinique d’un patient. Il peut s’agir de pathologies et symptômes, ou plus généralement de mots comme "entre", "rapporte" ou "continue". TIMEX3 d’expressions temporelles (une expression temporelle) Il s’agit d’expressions temporelles comme des dates, heures, durées, fréquences, ou intervalles. RML de résultats et mesures (un résultat ou une mesure) Il s’agit de résultats d’analyses de laboratoire, de mesures formelles, et de valeurs de mesure. ACTOR d’acteurs (un acteur) Il s’agit de patients, de professionnels de santé, ou d’autres acteurs pertinents pour l’histoire clinique d’un patient.
Table 8: Description of the NER tags used in our experiments for French.
Tag Tag name (in singular) Description PER nombres de personas (un nombre de persona) Estos son nombres de personas, ya sean reales o personajes ficticios. FAC instalaciones (una instalación) Estos son nombres de estructuras hechas por el hombre como infraestructura, edificios y monumentos. LOC lugares (un lugar) Estos son nombres de ubicaciones geográficas como hitos, ciudades, países y regiones. ORG organizaciones (una organización) Estos son nombres de organizaciones como empresas, agencias y partidos políticos. ACTI actividades y comportamientos (una actividad o comportamiento) Estas son palabras que se refieren a actividades humanas, comportamientos o eventos, así como actividades gubernamentales o regulatorias. ANAT anatomía (una anatomía) Estas son palabras que se refieren a la estructura del cuerpo humano, sus órganos y su posición, como partes del cuerpo u órganos, sistemas, tejidos, células, sustancias corporales y estructuras embrionarias. CHEM productos químicos y medicamentos (un producto químico o un medicamento) Estas son palabras que se refieren a una sustancia o composición que tiene una característica química, especialmente una propiedad curativa o preventiva con respecto a las enfermedades humanas o animales, como medicamentos, antibióticos, proteínas, hormonas, enzimas y sustancias peligrosas o venenosas. CONC conceptos e ideas (un concepto o una idea) Estas son palabras que se refieren a un concepto o una idea, como una clasificación, un producto intelectual, un idioma, una ley o un reglamento. DEVI dispositivos médicos (un dispositivo) Estas son palabras que se refieren a un dispositivo médico utilizado para administrar atención o realizar investigaciones médicas. DISO trastornos (un trastorno) Estas son palabras que se refieren a una alteración de la morfología, la función o la salud de un organismo vivo, animal o vegetal, como anomalías congénitas, disfunción, lesiones, signos o síntomas u observaciones. GENE genes y secuencias moleculares (un gen o una secuencia molecular) Estas son palabras que se refieren a un gen, un genoma o una secuencia molecular. GEOG áreas geográficas (un área geográfica) Estas son palabras que se refieren a un país, una región o una ciudad. LIVB seres vivos (un ser vivo) Estas son palabras que se refieren a un ser vivo o un grupo de seres vivos, como una persona o un grupo de personas, una planta o una categoría de plantas, un animal o una categoría de animales. OBJC objetos (un objeto) Estas son palabras que se refieren a cualquier cosa animada o inanimada que afecte los sentidos, como objetos físicos fabricados. OCCU ocupaciones (una ocupación) Estas son palabras que se refieren a una ocupación o disciplina profesional. ORGA organizaciones (una organización) Estas son palabras que se refieren a una organización, por ejemplo organizaciones relacionadas con la salud. PHEN fenómenos (un fenómeno) Estas son palabras que se refieren a un fenómeno que ocurre naturalmente o como resultado de una actividad, por ejemplo una función biológica.
Table 9: Description of the NER tags used in our experiments for Spanish.
Tag Tag name (in singular) Description PHYS fisiología (una fisiología) Estas son palabras que se refieren a cualquier elemento que contribuya al funcionamiento mecánico, físico y bioquímico o la organización de los organismos vivos y sus componentes. PROC procedimientos (un procedimiento) Estas son palabras que se refieren a una actividad o un procedimiento que contribuye al diagnóstico o tratamiento de pacientes, la información de pacientes, la capacitación del personal médico o la investigación biomédica. EVENT eventos (un evento) Estas son palabras que se refieren a acciones, estados y circunstancias que son relevantes para la historia clínica de un paciente, como patologías y síntomas, o más generalmente palabras como "entra", "reporta" o "continúa". TIMEX3 expresiones de tiempo (una expresión de tiempo) Estas son expresiones de tiempo como fechas, horas, duraciones, frecuencias o intervalos. RML resultados y mediciones (un resultado o una medida) Estos son resultados de análisis de laboratorio, mediciones formales y valores de medición. ACTOR actores (un actor) Estas son palabras que se refieren a pacientes, profesionales de la salud u otros actores relevantes para la historia clínica de un paciente. Abbreviation abreviaciones (una abreviación) Estas son los casos de siglas y acrónimos. Body_Part partes del cuerpo (una parte del cuerpo) Estas son palabras que se refieren a òrganos y partes anatómicas de personas. Clinical_Finding hallazgos clínicos (un hallazgo clínico) Estas son palabras que se refieren a observaciones, juicios o evaluaciones que se hacen sobre los pacientes. Diagnostic_Procedure procedimientos diagnósticos (un procedimiento diagnóstico) Estas son palabras que se refieren a exámenes que permiten determinar la condición del individuo. Disease enfermedades (una enfermedad) Estas son palabras que describen una alteración del estado fisiológico en una o varias partes del cuerpo, por causas en general conocidas, manifestada por síntomas y signos característicos, y cuya evolución es más o menos previsible. Family_Member miembros de la familia (un miembro de la familia) Estas son palabras que se refieren a miembros de la familia. Laboratory_or _Test_Result resultados de exámenes de laboratorio u otras pruebas (un resultado de un examen de laboratorio u otra prueba) Estas son palabras que se refieren a cualquier medición o evaluación obtenida a partir de un exámen de apoyo diagnóstico. Laboratory_Procedure procedimientos de laboratorio (un procedimiento de laboratorio) Estas son palabras que se refieren a exámenes que se realizan en diversas muestras de pacientes que permiten diagnosticar enfermedades mediante la detección de biomarcadores y otros parámetros. Se consideran los análisis de sangre, orina, y otros fluidos y tejidos que emplean métodos bioquímicos, microbiológicos y/o citológicos. Medication medicamentos o drogas (un medicamento o una droga) Estas son palabras que se refieren a medicamentos o drogas empleados en el tratamiento y/o prevención de enfermedades, incluyendo marcas comerciales y genéricos, así como también nombres para grupos de medicamentos. Procedure procedimientos (un procedimiento) Estas son palabras que se refieren a actividades derivadas de la atención y el cuidado de los pacientes. Sign_or_Symptom signos o síntomas (un signo o un síntoma) Estas son palabras que se refieren a manifestaciones de una enfermedad, determinadas mediante la exploración médica o percibidas y expresadas por el paciente. Therapeutic_Procedure procedimientos terapéuticos (un procedimiento terapéutico) Estas son palabras que se refieren a actividades o tratamientos que es empleado para prevenir, reparar, eliminar o curar la enfrmedad del individuo.
Table 10: Description of the NER tags used in our experiments for Spanish, continued.
A.4 Carbon footprint
Tables 11 and 12 detail the carbon emission estimations for all of our experiments. These estimations were computed with GreenAlgorithms v2.2 Lannelongue et al. (2021)  [^5] http://calculator.green-algorithms.org/, based on factors such as runtime, computing hardware and location where electricity used by our computer facility was produced. In total, the experiments described in this paper are estimated to have generated around 31kg of CO2 equivalent (29kg for the main experiments, and 2kg for ablation).
English French Spanish # Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL Few-shot approaches Causal 1 LLAMA-2-70B 46 44 126 233 54 85 131 129 273 284 41 76 114 344 2 LLAMA-3-8B-Instruct 3 5 12 19 12 5 7 7 11 23 4 7 12 36 3 Mistral-7B 4 6 12 24 8 5 8 14 13 25 7 5 11 27 4 Phi-3-medium-instruct 5 8 14 24 6 12 9 14 17 25 9 15 19 28 5 BLOOM-7B1 4 6 10 26 9 8 13 9 26 20 4 8 8 18 6 Falcon-40B 49 45 56 176 45 31 58 75 162 129 33 25 82 99 7 GPT-J-6B 7 6 8 23 7 5 8 13 21 17 6 6 13 28 8 OPT-66B 73 50 120 253 96 38 64 138 273 240 57 52 106 247 9 Vicuna-13B 10 11 20 52 11 11 12 18 33 40 10 11 22 51 10 Vicuna-7B 6 8 14 17 6 5 10 10 24 14 8 6 13 27 11 BioMistral-7B 7 8 11 17 5 7 12 10 28 14 8 8 11 23 12 Medalpaca-7B 8 4 17 24 10 7 14 11 19 13 7 8 15 26 13 Vigogne-13B 14 14 29 37 11 13 20 26 36 39 11 14 32 44 Masked 11 mBERT 2 1 2 2 2 2 2 2 1 1 1 2 1 2 12 XLM-R-large 2 2 2 1 2 2 2 2 2 2 1 1 1 2 13 BERT-large 2 1 2 2 2 - - - - - - - - - 14 RoBERTa-large 1 2 2 2 2 - - - - - - - - - 15 Bio_ClinicalBERT 2 2 1 2 1 - - - - - - - - - 16 ClinicalBERT 1 1 2 2 1 - - - - - - - - - 17 MedBERT 2 2 1 1 1 - - - - - - - - - 18 CamemBERT-large - - - - - 1 1 1 2 2 - - - - 19 FlauBERT-large - - - - - 2 2 2 2 2 - - - - 20 DrBERT-4GB - - - - - 2 2 2 2 2 - - - - 21 CamemBERT-bio - - - - - 1 2 2 2 2 - - - - 23 BETO - - - - - - - - - - 2 1 1 1 23 PatanaBERT - - - - - - - - - - 2 2 2 2 24 TulioBERT - - - - - - - - - - 1 2 2 1 25 BSC-BioEHR - - - - - - - - - - 2 2 2 2 26 BSC-Bio - - - - - - - - - - 2 2 2 2 Masked fully-supervised (skyline) RoBERTa-large 647 68 5 12 24 - - - - - - - - - CamemBERT-large - - - - - 595 15 4 5 8 - - - - BETO - - - - - - - - - - 579 41 3 21
Table 11: This table presents the carbon emissions (in g) of the optimization on the validation set of each model over each dataset. For CLMs, this correponds to the tree search over the prompt features through cross-validation. For MLMs, this corresponds to the supervised fine-tuning and training of the model.
English French Spanish # Model WikiNER CoNLL2003 E3C n2c2 NCBI WikiNER QFP E3C EMEA MEDLINE WikiNER CoNLL2002 E3C CWL Few-shot approaches Causal 1 LLAMA-2-70B 812 147 36 196 33 508 11 13 92 47 514 201 11 198 2 LLAMA-3-8B-Instruct 240 39 9 63 22 158 3 4 29 21 279 53 2 34 3 Mistral-7B 234 35 8 59 21 148 3 4 27 20 261 50 2 32 4 Phi-3-medium-instruct 326 48 12 64 25 380 5 9 62 48 529 70 4 76 5 BLOOM-7B1 220 33 8 44 16 255 3 5 38 29 261 47 2 46 6 Falcon-40B 600 109 26 144 46 722 9 19 155 70 752 154 9 157 7 GPT-J-6B 146 17 4 53 20 245 2 6 14 26 154 40 3 53 8 OPT-66B 765 139 33 185 63 971 12 27 179 93 993 196 12 217 9 Vicuna-13B 314 47 11 63 24 363 5 8 61 46 502 67 4 74 10 Vicuna-7B 146 17 4 53 20 246 2 6 14 26 155 65 3 53 11 BioMistral-7B 235 35 9 49 17 269 3 5 43 32 272 49 2 48 12 Medalpaca-7B 192 24 5 39 14 98 2 2 17 13 172 53 1 21 13 Vigogne-13B 322 49 11 65 24 245 5 6 44 33 361 68 3 66 Masked 11 mBERT 14 4 <1 2 <1 15 1 <1 1 1 13 2 <1 2 12 XLM-R-large 14 4 <1 2 <1 15 1 <1 1 1 13 2 <1 2 13 BERT-large 14 4 <1 2 <1 - - - - - - - - - 14 RoBERTa-large 14 4 <1 2 <1 - - - - - - - - - 15 Bio_ClinicalBERT 14 4 <1 2 <1 - - - - - - - - - 16 ClinicalBERT 14 4 <1 2 <1 - - - - - - - - - 17 MedBERT 14 4 <1 2 <1 - - - - - - - - - 18 CamemBERT-large - - - - - 15 1 <1 1 1 - - - - 19 FlauBERT-large - - - - - 15 1 <1 1 1 - - - - 20 DrBERT-4GB - - - - - 17 1 <1 1 1 - - - - 21 CamemBERT-bio - - - - - 15 1 <1 1 1 - - - - 22 BETO - - - - - - - - - - 13 2 <1 2 23 PatanaBERT - - - - - - - - - - 13 2 <1 2 24 TulioBERT - - - - - - - - - - 13 2 <1 2 25 BSC-BioEHR - - - - - - - - - - 13 2 <1 2 26 BSC-Bio - - - - - - - - - - 13 2 <1 2 Masked fully-supervised (skyline) RoBERTa-large 14 4 <1 2 <1 - - - - - - - - - CamemBERT-large - - - - - 15 1 <1 1 1 - - - - BETO - - - - - - - - - - 13 2 <1 2
Table 12: This table presents the carbon emissions (in g) of the inference on the test set of each model over each dataset.